Grade 12th passMechanics

A particle of mass m is projected with velocity v making of an angle of 45 degree with the
horizontal. When the particle lands on the level ground, then what is the magnitude of the
change in its momentum?

Preeti
10 Years agoGrade 12th pass

1 Answer

Riddhish Bhalodia
10 Years ago
As there is no acceleration in the x direction, the velocity along x direction remains constant and hence no change in momentum along x axis
at the time of landing the velocity of particle in y axis is exactly opposite to that of the time when the particle was thrown.
Hence change in velocity = 2v/\sqrt{2} = v\sqrt{2}
thus change in momentum in magnitude is = mv\sqrt{2}